OpenStudy (anonymous):

You have 29 coins (only quarters and dimes). They are worth $4.55, show work using the elimination method to find out how many quarters you have.

OpenStudy (johnweldon1993):

well you have 2 equations here q + d = 29 (quarters + dimes = total coins) .25q + .10d = 4.55 (worth * amount + worth * amount = total amount) So you can use the elimination method from here....can you do this?

OpenStudy (anonymous):

I'm confused on how the elimination method works. Also where did you get the 10 from?

OpenStudy (johnweldon1993):

.10....is the worth of the dime....and .25 is the worth of the quarter. right? a dime is 10 cents and a quarter is 25

OpenStudy (johnweldon1993):

and the elimination method works by eliminating 1 variable so you may solve for the other....I'll walk you through it if you would like...

OpenStudy (anonymous):

Yes that would be great :)

OpenStudy (johnweldon1993):

so you have your 2 equations q + d = 29 .25q + .10d = 4.55 What you want to do is multiply either the top or bottom equation...by a value...that will make it equal with the value in the equation below it...doing this...will eliminate a variable....in this case q + d = 29 .25q + .10d = 4.55 notice how if you multiply the TOP equation...by .10 you'll get .10q + .10d = 2.9 now this with the second equation below it you have .10q + .10d = 2.9 .25q + .10d = 4.55 Now what you want to do...is subtract these 2 equations...why? because as you can see....the value of 'd' is the same in the top equation....and the bottom equation...so subtracting them will eliminate 'd'

OpenStudy (johnweldon1993):

so when you do that .10q + .10d = 2.9 -(.25q + .10d = 4.55) you get -.15q + 0 = -1.65 (0 because there are no more 'd' left Now you ONLY have 1 variable....'q' that is what you are going to solve for first here divide both sides by -.15 -.15q = -1.65 ------------- -.15 -.15 and you get 11 This is how many quarters you have To find out how many dimes....you just look at your first ORIGINAL equation....and put 11 in for 'q' q + d = 29 we know q = 11 so 11 + d = 29 -11 -11 d = 18 Now you know you have 11 quarters...and 18 dimes

OpenStudy (johnweldon1993):

Hope that's understandable! :)

OpenStudy (anonymous):

Wow, that actually was very helpful, and easy to understand! Thanks a lot :D

OpenStudy (johnweldon1993):

No problem! :)